1. c) The radius of a silver bar is 30 cm. What force is required to stretch the bar by 12.7% of its length where Young's modulus is Y = 7.25 * 10 ^ 10 * N * m ^ - 2 .

1
Expert's answer
2021-12-06T16:43:04-0500
"7.25\cdot10^{10}=\frac{F}{\pi(0.3)^20.127}\\F=2.60\cdot10^{9}\ N"